Overview DC/EC - Standard font set

The following table describes the generated *.mf files which are needed for the setup of the standard DC/EC-font set.

Fonts  5  6  7  8  9 10 10.95[11] 12 14.4[14] 17.28[17]
dccr  •  •  •  •  •  •  •  •  •  •